OpenStudy (anonymous):

What is the equation of a circle with center (2, -5) and radius 4?

OpenStudy (anonymous):

The equation of a circle is (x-k)^2+(x-h)^2= r^2 k= the x value you are given in the ordered pair h= the y value you are given in the ordered pair so plug it in and do (x-2)^2+(x-(-5))^2= 4^2 then you will get (x-2)^2+ (x+5)^2= 16 and there you have your answer.

OpenStudy (anonymous):

Thank you!